# The Generation Phases of Flow Matching: a Denoising Perspective

### v1

@@ Appendix F Perturbation samples
Figure 11: Unperturbed images.
@@
Appendix G Samples obtained on CelebA-64
(a) $\mathcal{L}_{\text{den}},\mathcal{C}_{\text{NN}}$
(b) $\mathcal{L}_{\text{FM}},\mathcal{C}_{\text{NN}}$
(c) $\mathcal{L}_{\text{den}},\mathcal{C}_{\mathrm{I+NN}}$
(d) $\mathcal{L}_{\text{FM}},\mathcal{C}_{\mathrm{I+NN}}$
(e) $\mathcal{L}_{\text{mid-0.2}},\mathcal{C}_{\mathrm{NN}}$
(f) $\mathcal{L}_{\text{mid}},\mathcal{C}_{\text{NN}}$
(g) $\mathcal{L}_{\text{den}},\mathcal{C}_{\text{NN}}$
(h) $\mathcal{L}_{\text{FM}},\mathcal{C}_{\text{NN}}$
(i) $\mathcal{L}_{\text{den}},\mathcal{C}_{\mathrm{I+NN}}$
(j) $\mathcal{L}_{\text{FM}},\mathcal{C}_{\mathrm{I+NN}}$
(k) $\mathcal{L}_{\text{mid-0.2}},\mathcal{C}_{\mathrm{NN}}$
(l) $\mathcal{L}_{\text{mid}},\mathcal{C}_{\text{NN}}$
Figure 13: Sampling results on CelebA-64 for different models and two initial points $x_{0}$ (shared across columns).
@@ H.1 Time-interval Jacobian penalized models
We train standard flow matching models ($\mathcal{L}_{\text{FM}},\mathcal{C}_{\mathrm{I+NN}}$) with additional Jacobian spectral norm regularization applied over a prescribed time interval $t_{\mathrm{min}},t_{\mathrm{max}}]$:
@@
where $\lambda$ is a regularization parameter and $M$ is the targeted upper bound on the Jacobian spectral norm. The models are trained without regularization for $390$ epochs and finetuned with regularization for the last $10$ epochs. The Jacobian spectral norm is estimated using the power method ($10$ iterations). The threshold $M$ is set to $M=4$ for early intervals (i.e. $[0,0.2],[0.1,0.3]$) and $M=2$ for the others. The regularization parameter is set to $\lambda=0.1$ for all intervals except for $[0.3,0.6]$ and $[0.5,0.8]$ where a more aggressive regularization $\lambda=0.2$ is used to ensure an effective reduction of the Jacobian spectral norm on these intervals. Indeed, our goal here is to induce a controlled decrease of the Jacobian norm over selected time intervals and study how this affects denoising and generation.
@@
The early-time regularizations (i.e. $[0,0.2]$ and $[0.1,0.3]$) produce the largest deviation from the closed-form behaviour in terms of the mean Jacobian spectral norm measured over the ODE trajectories (see Figure 14(a)): models with these early-time regularizations fail to reproduce the sharp peak of the closed-form Jacobian around $t\approx 0.2$ and instead maintain higher values in mid times. In contrast, models regularized at mid or late times begin by increasing the Jacobian spectral norm (similarly to standard FM) and then exhibit a decay around mid times, getting closer to the closed-form behaviour.
@@
In terms of denoising performance (Figure 15(a)), a decrease of the Jacobian norm over a given interval (compared to the standard FM model) corresponds to a decrease in PSNR at time steps within that interval. In other words, constraining the local Lipschitz constant deteriorates the denoising accuracy at these times.
In terms of generation performance (Figure 15(b)), two classes of models clearly appear. Penalizing the Jacobian norm in early intervals ($[0,0.2]$, $[0.1,0.3]$) achieves FID scores comparable to standard FM. In contrast, penalizing it in the other intervals results in higher FIDs (especially for mid time intervals $[0.3,0.6]$, $[0.3,0.8]$). To better quantify the impact of these regularizations on generated samples, we compute the average pairwise distance between samples generated from the same $x_{0}$, thus directly comparing ODE trajectories (Figure 17(a)). Among all models, larger distances relative to the standard FM baseline are observed for penalizations applied at early and mid times. Qualitatively (see Figure 18), early-time penalization ($[0.1,0.3]$) can lead to visually distinct samples, sometimes even changing the image class, whereas late-time regularization mostly yields small perturbations or noisy versions compared to the standard FM samples.
This aligns with observations made from the controlled perturbation experiments in Section 6.1, where macro-level perturbations applied at early times tend to shift the global image distribution but maintain a low FID, whereas micro-level perturbations applied at late times produce noisier samples and lead to a stronger degradation of the FID.
@@ Discussion
Overall, this experiment indicates that shifting the peak of the Jacobian spectral norm as done with the early-time penalizations can change the ODE trajectories without altering the generation quality. In contrast, maintaining a relatively high Jacobian norm appears important in order to keep good denoising quality in the mid times, which in turn is crucial to achieve good generation quality. Notably, matching the behaviour of the closed-form velocity field in terms of the Jacobian spectral norm seems undesirable in the early/mid-time regimes: not reproducing its peak at early times still yields models with competitive FID scores, whereas getting closer to the closed-form behaviour at mid times results in degraded FIDs. This can be understood from the perspective of the closed-form denoiser, which at mid times reproduces samples from the training dataset: distinct input points are mapped to the same output, i.e. image of the training set, which is not the expected behaviour of a good denoiser or a good generator.
@@ H.2 Models with intermediate weighting
Surprisingly, despite the formal resemblance between the standard flow-matching weighting $w_{t}^{\mathrm{FM}}=\frac{1}{1-t}^{2}$ and the limiting behavior of $w_{t}^{\mathrm{mid-}t^{*}}$ as $t^{*}\to 1$, the results diverge markedly. Instead of recovering the performance of the FM baseline, we observe a sharp degradation: the FID explodes for $t^{*}=0.95$, and the PSNR deteriorates across all times, including $t=0.95$ itself. This finding indicates that it is not straightforward to isolate the influence of a single time point: poor denoising performance at other times propagates through the training dynamics, affecting even the regions of emphasis. Interestingly, the model most similar to the FM-shifted baseline is obtained for $t^{*}=0.2$. These effects are not attributable to the architectural choice of class $\mathcal{C}_{\mathrm{NN}}$ versus $\mathcal{C}_{\mathrm{I+NN}}$, since the pair $(\mathcal{L}_{\mathrm{FM}},\mathcal{C}_{\mathrm{NN}})$ also has a very good FID.
Two distinct degradation regimes emerge in the FID/PSNR trends, echoing the observations of Section 6.1. In the first phase, when the emphasis is placed on intermediate times $t^{*}\in[0.2,0.5]$, the FID remains roughly constant and close to that of the uniform weighting $w_{t}^{\mathrm{den}}=1$, yet the samples display visible drifts (Figure 19). These models appear to have learned qualitatively different denoising functions, offering insights into how weighting can modulate trajectory alignment. In the second phase, corresponding to larger times $t^{*}\in[0.6,0.9]$, image quality degrades sharply, both in terms of FID and PSNR.
Pairwise distance analysis corroborates these observations: the distance from samples generated using the baseline model increases monotonically with $t^{*}$, with a mild peak around $t^{*}=0.7$. We hypothesize that this pattern reflects two regimes: a “drift perturbation” regime for lower $t^{*}$, where structural shifts dominate, and a “noise amplification” regime for higher $t^{*}$, where instability in late-time dynamics prevails. Indeed, models with $t^{*}=0.5$ and $t^{*}=0.8$ exhibit comparable distances to the baseline while producing qualitatively distinct images (Figure 19).
Finally, these findings align with our analysis of the Lipschitz constant. Models that drastically reduce their Lipschitz constant relative to the FM baseline in the interval $t\in[0.3,0.6]$ coincide with those exhibiting catastrophic FID degradation. Conversely, “drift-perturbed” models tend to suppress the Lipschitz constant primarily for $t\in[0,0.3]$, while maintaining higher expressivity in the mid-range $t\in[0.3,0.6]$.
@@
Figure 14: Mean and standard deviation of the spectral norm $\|\nabla_{x}v(x_{t},t)\|_{2}$, computed on 1000 ODE trajectories $(x_{t})$. CIFAR-10, 400 epochs. The Jacobian spectral norm is estimated using the power iteration with maximum number of iteration set to $10$.
@@
(b) FID on 10k tests images.
Figure 15: PSNR and FID for the different regularizations, CIFAR-10, 400 epochs. PSNR degradation at early times does not impair generation performance, while PSNR degradation at mid times systematically correlates with a higher FIDs.
@@
(b) FID on 10k tests images.
Figure 16: PSNR and FID for the different mid losses, CIFAR-10, 400 epochs.
@@
Figure 17: Average pairwise distance inter-models computed on $1000$ samples, sharing same $x_{0}$ across models. Experiments done on CIFAR-10, models trained with 400 epochs.
$\mathcal{L}_{\mathrm{FM}},\ \mathcal{C}_{\mathrm{I+NN}}$ $\mathcal{L}_{\mathrm{FM}}+\mathcal{R}_{0.1-0.3},\ \mathcal{C}_{\mathrm{I+NN}}$ $\mathcal{L}_{\mathrm{FM}}+\mathcal{R}_{0.5-0.8},\ \mathcal{C}_{\mathrm{I+NN}}$
@@
Figure 18: CIFAR-10 samples generated by models trained with different Jacobian regularizations. Regularizing at early times ($\mathcal{R}_{0.1-0.3}$) changes visually more the samples than applying regularization at later times ($\mathcal{R}_{0.5-0.8}$).
$\mathcal{L}_{\mathrm{FM}},\ \mathcal{C}_{\mathrm{I+NN}}$ $\mathcal{L}_{\mathrm{mid0.5}},\ \mathcal{C}_{\mathrm{NN}}$ $\mathcal{L}_{\mathrm{mid0.8}},\ \mathcal{C}_{\mathrm{NN}}$
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption

### v2

@@ Appendix G Ablation on the perturbation experiments Section 5
Figure 13: Frequency-based perturbations applied to denoiser $D_{t}$ (here for $t=0.3)$. Experiments done here on CelebA 128.
Refer to caption Refer to caption Refer to caption
FM baseline sample 1 FM baseline sample 2 FM baseline sample 3
Interval Ckb. 4x4 Ckb. 8x8 Ckb. 16x16 Ckb. 32x32 Ckb. 64x64 Pos. shift Neg. shift
$[0.0,0.3]$ Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
$[0.0,0.3]$ Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
$[0.0,0.3]$ Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Figure 14: Effect of perturbations on generated CelebA $128\times 128$ samples.
Refer to caption Refer to caption Refer to caption
FM baseline sample 1 FM baseline sample 2 FM baseline sample 3
Interval Low freq. - 0.05 Low freq - 0.15 Low freq - 0.5 Low freq lum - 0.05 Low freq lum - 0.15 Low freq lum - 0.5
$[0.0,0.3]$ Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
$[0.3,0.6]$ Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
$[0.6,0.9]$ Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
$[0.0,0.3]$ Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
$[0.3,0.6]$ Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
$[0.6,0.9]$ Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
$[0.0,0.3]$ Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
$[0.3,0.6]$ Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
$[0.6,0.9]$ Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Figure 15: Effect of frequency-domain perturbations on CelebA $128\times 128$ samples.
Figure 16: Influence of different perturbations at different generation phases on the FID (10K, test) on CIFAR10.
Figure 17: Influence of different perturbations at different generation phases on the FID (10K, test) on CIFAR10.
Refer to caption Refer to caption
FM baseline sample 1 FM baseline sample 2
@@ H.1 Time-interval Jacobian penalized models
Our goal here is to induce a controlled decrease of the Jacobian norm over selected time intervals and study how this affects denoising and generation. We train standard flow matching models ($\mathcal{L}_{\text{FM}},\mathcal{C}_{\mathrm{I+NN}}$) with additional Jacobian spectral norm regularization applied over a prescribed time interval $t_{\mathrm{min}},t_{\mathrm{max}}]$:
@@
where $\lambda$ is a regularization parameter and $M$ is the targeted upper bound on the Jacobian spectral norm. The models are trained without regularization for $990$ epochs and finetuned with regularization for the last $10$ epochs. The Jacobian spectral norm is estimated using the power method ($10$ iterations). We choose the early phase interval $[0.1,0.3]$ (see Figure 7(a)) and three intermediate phase intervals, $[0.3,0.6]$, $[0.6,0.9]$ and $[0.4,0.8]$. The threshold $M$ is set to $M=4$ for $[0.1,0.3]$, and $M=2$ for $[0.4,0.8]$ and $[0.3,0.6]$, and $M=1$ for the $[0.6,0.9]$. The regularization parameter is set to $\lambda=0.1$ for $[0.1,0.3]$, $\lambda=0.2$ for $[0.3,0.6]$ and $[0.4,0.8]$, $\lambda=0.4$ for $[0.6,0.9]$.
@@
The early-time regularization (i.e. $[0.1,0.3]$) produces the largest deviation from the closed-form behaviour in terms of the mean Jacobian spectral norm measured over the ODE trajectories (see Figure 19(a)): the regularized model don’t reproduce the sharp peak of the closed-form Jacobian around $t\approx 0.2$ and instead maintain higher values in mid times. In contrast, models regularized in the intermediate time phase ($[0.3,0.6]$, $[0.6,0.9]$ and $[0.4,0.8]$) exhibit a decay of the Lipschitz constant around time 0.20.30.2-0.3 getting closer to the closed-form behaviour.
@@
In terms of denoising performance (Figure 20(a)), a decrease in the Jacobian norm over a given time interval (compared to the standard FM model) corresponds to a decrease in PSNR at the time steps within that interval. In other words, constraining the local Lipschitz constant degrades the denoising accuracy at those times.
In terms of generation performance (Figure 20(b)), two distinct behaviors emerge. Penalizing the Jacobian norm during the early phase yields FID scores comparable to, and slightly better than, the standard FM model. Qualitatively (see Figure 23), this produces samples that differ noticeably from the FM baseline, sometimes even altering the image class. In contrast, penalizing the model during the intermediate phase leads to significantly higher FIDs and noisier samples compared to the standard FM outputs. This is consistent with our controlled perturbation experiments in Section 5, where perturbations applied at early times tend to shift the global image distribution while maintaining a low FID, whereas perturbations applied at later times produce noisier samples and result in a more severe degradation of the FID.
To further quantify the impact of these regularizations on the generated samples, we compute the average pairwise distance between samples generated from the same $x_{0}$, thereby directly comparing their ODE trajectories (Figure 22(a)). Among all models, the largest distances (relative to the standard FM baseline) are observed for models regularized during the early phase or at the beginning of the intermediate phase.
@@ Discussion
Overall, this experiment indicates that shifting the peak of the Jacobian spectral norm as done with the early-time penalizations can change the ODE trajectories without altering the generation quality. In contrast, maintaining a relatively high Jacobian norm appears important in order to keep good denoising quality in the mid times, which in turn is crucial to achieve good generation quality. Notably, matching the behaviour of the closed-form velocity field in terms of the Jacobian spectral norm seems undesirable in the early/mid-time regimes: not reproducing its peak at early times still yields models with competitive FID scores, whereas getting closer to the closed-form behaviour at mid times results in degraded FIDs.
@@ H.2 Models with intermediate weighting
Surprisingly, despite the formal resemblance between the standard flow-matching weighting $w_{t}^{\mathrm{FM}}=\frac{1}{1-t}^{2}$ and the limiting behavior of $w_{t}^{\mathrm{mid-}t^{*}}$ as $t^{*}\to 1$, the results diverge markedly. Instead of recovering the performance of the FM baseline, we observe a sharp degradation: the FID explodes for $t^{*}=0.95$, and the PSNR deteriorates across all times. This finding indicates that it is not straightforward to isolate the influence of a single time point: poor denoising performance at other times propagates through the training dynamics, affecting even the regions of emphasis. Interestingly, the model most similar to the standard FM baseline is obtained for $t^{*}=0.2$.
Two distinct degradation regimes emerge in the FID/PSNR trends, consistent with the observations of Section 5. In the first regime, where the emphasis is placed on intermediate times $t^{\star}\in[0.2,0.5]$, the FID remains nearly unchanged and close to the SOTA baseline. Pairwise distances (Figure 22(b)) and PSNR curves (Figure 21(a)) show that, while denoising performance in the early phase worsens as $t^{\star}$ increases, this mainly induces sample drift rather than a loss in overall generation quality (see samples in Figure 24). These models appear to have learned qualitatively different denoising functions, highlighting how weighting choices can modulate trajectory alignment.
In the second regime, for larger times $t^{\star}\in[0.6,0.9]$, image quality degrades sharply: both FID and intermediate-time PSNR deteriorate, and pairwise distances decrease again for $t^{\star}>0.8$.
These observations suggest that while poor denoising performance in the early phase (e.g. weighting with $t^{\star}=0.5$) primarily induces sample drifts, poor denoising performance in the the intermediate phase (e.g. weighting with $t^{\star}=0.9$) mainly affect the generation quality.
@@
Figure 19: Mean and standard deviation of the spectral norm $\|\nabla_{x}v(x_{t},t)\|_{2}$, computed on 1000 ODE trajectories $(x_{t})$. CIFAR-10, 1000 epochs. The Jacobian spectral norm is estimated using the power iteration with maximum number of iteration set to $10$.
@@
(b) FID on 50k train images.
Figure 20: PSNR and FID for the different regularizations, CIFAR-10, 1000 epochs. PSNR degradation at early times does not impair generation performance, while PSNR degradation at mid times systematically correlates with a higher FIDs.
@@
(b) FID on 50k train images.
Figure 21: PSNR and FID for the different mid losses, CIFAR-10, 1000 epochs.
@@
Figure 22: Average pairwise distance inter-models computed on $1000$ samples, sharing same $x_{0}$ across models. Experiments done on CIFAR-10, models trained with 1000 epochs.
$\mathcal{L}_{\mathrm{FM}},\ \mathcal{C}_{\mathrm{I+NN}}$ $\mathcal{L}_{\mathrm{FM}}+\mathcal{R}_{0.1-0.3},\ \mathcal{C}_{\mathrm{I+NN}}$ $\mathcal{L}_{\mathrm{FM}}+\mathcal{R}_{0.4-0.8},\ \mathcal{C}_{\mathrm{I+NN}}$
@@
Figure 23: CIFAR-10 samples generated by models trained with different Jacobian regularizations. Regularizing at early times ($\mathcal{R}_{0.1-0.3}$) changes visually more the samples than applying regularization at intermediate times ($\mathcal{R}_{0.4-0.8}$).
$\mathcal{L}_{\mathrm{FM}},\ \mathcal{C}_{\mathrm{I+NN}}$ $\mathcal{L}_{\mathrm{mid0.2}},\ \mathcal{C}_{\mathrm{I+NN}}$ $\mathcal{L}_{\mathrm{mid0.5}},\ \mathcal{C}_{\mathrm{NN}}$ $\mathcal{L}_{\mathrm{mid0.9}},\ \mathcal{C}_{\mathrm{NN}}$
Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption
@@
$\mathcal{L}_{\mathrm{FM}},\ \mathcal{C}_{\mathrm{I+NN}}$ $\mathcal{L}_{\mathrm{mid0.2}},\ \mathcal{C}_{\mathrm{I+NN}}$ $\mathcal{L}_{\mathrm{mid0.5}},\ \mathcal{C}_{\mathrm{NN}}$ $\mathcal{L}_{\mathrm{mid0.8}},\ \mathcal{C}_{\mathrm{NN}}$
Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption
Figure 25: CelebA $128\times 128$ samples generated by models trained with different mid weightings. Each row corresponds to one fixed initial point $x_{0}$.
Appendix I Ablation on the ODE solver
Here, we perform a sanity check to ensure that the choice of ODE solver does not affect the relative performance of the models. In Table 4, we compare the FID scores of six models when samples are generated using dopri5 or euler100. The results show that the ranking of the models is consistent across both solvers, indicating that our choice of dopri5 does not bias the comparison.
Euler (100 steps) Dorpi5
$\mathcal{L}_{\text{FM}}$ $\mathcal{C}_{\mathrm{NN}}$ 5.93 (2) 4.89 (2)
$\mathcal{L}_{\text{FM}}$ $\mathcal{C}_{\mathrm{I+NN}}$ 4.84 (1) 3.90 (1)
$\mathcal{L}_{\text{den}}$ $\mathcal{C}_{\mathrm{NN}}$ 13.37 (4) 16.02 (4)
$\mathcal{L}_{\text{den}}$ $\mathcal{C}_{\mathrm{I+NN}}$ 10.51 (3) 10.29 (3)
$\mathcal{L}_{\text{classic}}$ $\mathcal{C}_{\mathrm{NN}}$ 54.09 (5) 71.07 (5)
$\mathcal{L}_{\text{classic}}$ $\mathcal{C}_{\mathrm{I+NN}}$ 113.92 (6) 116.18 (6)
Table 4: FID 50k on train set. The numbers in blue are the ranking of the models.
Appendix J Justification of the FM weighting
In the following, we attempt to provide a statistical interpretation for why the standard Flow Matching weighting
$w(t)=\frac{1}{(1-t)^{2}}$
often yields the best empirical performance. This is not meant to be a formal proof, but rather a conceptual link with classical results in the regression literature [shalizi2013weighted, schick1997efficient, aitken1936iv].
Maximum-likelihood justification on a toy model.
We consider the following toy generative model. Let
$X_{0}\sim\mathcal{N}(0,I),\qquad X_{1}\sim\mathcal{N}(0,\tau^{2}I),\qquad X_{0}\perp X_{1},$
and for $t\in(0,1)$ define $X_{t}=(1-t)X_{0}+tX_{1}$. Introducing $Y=X_{t}/t$ yields
$Y=X_{1}+\sigma(t)X_{0},\qquad\sigma(t)=\frac{1-t}{t},$
so that
$Y\mid(X_{1},t)\sim\mathcal{N}(X_{1},\sigma^{2}(t)I).$
In this linear–Gaussian setting, the posterior distribution admits a closed form:
$X_{1}\mid(Y=y,t)\sim\mathcal{N}\bigl(\mu_{\mathrm{post}}(y,t),\,\sigma_{\mathrm{post}}(t)I\bigr),$
with
$\sigma_{\mathrm{post}}(t)=\left(\frac{1}{\tau^{2}}+\frac{1}{\sigma^{2}(t)}\right)^{-1},\qquad\mu_{\mathrm{post}}(y,t)=\frac{\tau^{2}}{\tau^{2}+\sigma^{2}(t)}\,y.$
Since our goal is to approximate the conditional expectation $\mathbb{E}[X_{1}\mid X_{t}=x_{t},t]$ (or equivalently $\mathbb{E}[X_{1}\mid Y=y,t]$), we introduce a parametric function $f_{\theta}$ intended to model this conditional mean. We then define a conditional Gaussian model
$p_{\theta}(x_{1}\mid y,t)=\mathcal{N}\bigl(x_{1};f_{\theta}(y,t),\sigma^{2}_{\mathrm{post}}(t)I\bigr).$
Of course, in this toy example we already know that $\mathbb{E}[X_{1}\mid Y=y,t]=\mu_{\mathrm{post}}$ but we still want to estimate it for the sake of understanding the weightings. For a single sample the negative log-likelihood (NLL) is
$\log p_{\theta}(x_{1}\mid y,t)=\frac{1}{2\sigma^{2}_{\mathrm{post}}(t)}\|x_{1}-f_{\theta}(y,t)\|^{2}+\frac{d}{2}\log\bigl(2\pi\sigma^{2}_{\mathrm{post}}(t)\bigr),$
so the expected NLL is, up to an additive constant,
$\mathcal{L}_{\mathrm{NLL}}(\theta)=\mathbb{E}\Bigl[\frac{1}{\sigma^{2}_{\mathrm{post}}(t)}\,\|X_{1}-f_{\theta}(Y,t)\|^{2}\Bigr].$
Hence a time-dependent weighting appears naturally:
$w(t)\propto\frac{1}{\sigma^{2}_{\mathrm{post}}(t)}=\frac{1}{\tau^{2}}+\frac{t^{2}}{(1-t)^{2}}.$
This weight behaves as $1/(1-t)^{2}$ for large $t$. The toy model therefore provides a principled statistical motivation: inverse-variance weighting based on the forward corruption noise $\sigma^{2}(t)$ is the maximum-likelihood choice in an analytically tractable setting, and this coincides with the empirically best-performing weighting $(1/(1-t))^{2}$.
Remark 2.
We also experiment with an SNR-based weighting, given by $w_{t}=\frac{t^{2}}{(1-t)^{2}}$, which corresponds to the signal-to-noise ratio of $x_{t}$. We report the corresponding performance in Table 2. We find that this weighting performs similarly to the weighting $1/(1-t^{2})$, and even slightly better. This is also consistent with empirical practices in diffusion models, where a similar weighting has been used (see the discussion in the related works, Appendix A).
Remark 3.
If we do not assume $X_{1}$ to have a Gaussian prior, one may instead adopt a different modeling assumption and approximate the posterior $p(x_{1}\mid y,t)$ by a Gaussian distribution with mean $\mathbb{E}[X_{1}\mid Y=y,t]$ and covariance $C_{t}:=\mathrm{Cov}[X_{1}\mid Y=y,t]$. Under this assumption, the same maximum-likelihood reasoning leads to the weighted quadratic loss
$\mathbb{E}\Bigl[(X_{1}-f_{\theta}(Y,t))^{\top}C_{t}^{-1}(X_{1}-f_{\theta}(Y,t))\Bigr].$
This loss is however intractable in practice since $C_{t}$ depends on the unknown data distribution.
Although this Gaussian approximation is not valid in general, it becomes increasingly reasonable near $t=1$, where the conditional distribution $X_{1}\mid Y=y,t$ concentrates around a single point. Moreover, we expect $C_{t}\to 0$ as $t\to 1$, which justifies the use of a weighting $w_{t}$ that diverges near $t=1$.
Connection with heteroscedastic regression in the linear case.
We now assume that the estimator $f$ is restricted to a linear form. Consider the regression model
$X_{1}=Y\beta+\varepsilon(t),$
where $A$ is a linear operator and $\varepsilon$ is a zero-mean Gaussian noise satisfying
$\mathbb{E}[\varepsilon\mid Y,t]=0,\qquad\mathrm{Cov}[\varepsilon(t)\mid Y,t]=v(t)I.$
The model is said to be heteroscedastic because the conditional covariance $v(t)$ is not constant but varies with $t$.
The Generalized least square result [aitken1936iv, greene2003econometric] states that, the best linear unbiased estimator (referred to as BLUE) of $A$, i.e. the one with minimum covariance (in the sense of positive semidefinite ordering) is obtained by solving the generalized least-squares problem
$\hat{A}=\arg\min_{A}\mathbb{E}\left[\frac{1}{v(t)}\bigl\|X_{1}-AY\bigr\|^{2}\right].$
Thus, in the linear case, the optimal weighting is proportional to the inverse conditional variance of the noise, $\frac{1}{v(t)}$. We can therefore see the case of Flow Matching as a generalization of this principle.
Appendix K Faster training with the residual parametrization
Here we provide empirical evidence that the residual parametrization $\mathcal{C}_{\mathrm{I+NN}}$ not only achieves better final performance, but also converges faster, requiring fewer training epochs to reach a given FID than the $\mathcal{C}_{\mathrm{NN}}$ parametrization. In Figure 26, we report the FID (10k test) on CelebA $128\times 128$ as a function of the number of training epochs.
Figure 26: FID (10k test) on CelebA $128\times 128$ as a function of training epochs.
Appendix L Additional experiment on the regularity of target and learned models
We provide in Figure 27 additional evidence of the gap between the regularity of a trained model (here, the standard flow matching baseline corresponding to loss $\mathcal{L}_{\mathrm{FM}}$ and parameterization $\mathcal{C}_{\mathrm{I+NN}}$) and the closed-form target. In this experiment, we use the TinyImagenet dataset, which is a downscaled variant of the ImageNet dataset that contains 200 classes with 500 training images per class and image resolution $\times64$, using the training code and the backbone architecture provided in lipman2024flowmatchingguidecode.
Figure 27: Mean and standard deviation of the spectral norm $\|\nabla_{x}v(x_{t},t)\|_{2}$ computed on ODE trajectories ($1000$ points for the closed-form velocity, $400$ for the standard FM trained model $\mathcal{L}_{\mathrm{FM}},\mathcal{C}_{\mathrm{I+NN}}$) on TinyImagenet with Gaussian $p_{0}$.